# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v1

@@ III FAJSCC Framework
Our overall FAJSCC structure is illustrated in Figure 2. To facilitate deepJSCC training, standardization of the encoder maps pixel values of an image from $[0,255]$ to $[-1,1]$ while destandardization of the decoder reverses this process. Power normalization of the encoder rescales features to fit power constraint $P=1$. Following previous works [41, 16, 8, 10], our FAJSCC uses patch embedding to refine the image at the feature level, patch merging for down-sampling, and patch division for up-sampling. Based on the feature sizes, both the encoder and decoder consist of multiple stages, indexed by $i\in[1:L]$, where $L$ is the total number of stages. The patch merging layer of the $i$-th stage transforms the feature dimension from $\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times C_{i-1}$ to $\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{i}$. Conversely, the patch division layer in the $i$-th stage transforms the feature dimensions from $\frac{H}{2^{L+1-i}}\times\frac{W}{2^{L+1-i}}\times C_{L+1-i}$ to $\frac{H}{2^{L-i}}\times\frac{W}{2^{L-i}}\times C_{L-i}$. Here, $C_{0}$ is the number of RGB color channels, i.e., $C_{0}=3$. The encoder’s fully connected (FC) layer scales input feature channel dimensions to fit the available bandwidth while the decoder’s FC layer adjusts input feature channel dimensions to align the feature channel sizes of the decoder’s first stage.
In each stage, the feature importance-aware group (FA group), consisting of multiple feature importance-aware blocks (FA blocks), performs the core encoding and decoding processes of FAJSCC. To enhance communication accuracy while reducing computational resources, our FA block employs importance-wise feature processing and unified lightweight computation. Importance-wise feature processing prioritizes computational resources for important features to increase computational efficiency. Unified lightweight computation leverages feature correlations for error correction while filtering out unimportant or noisy information, ensuring reliable image transmission and reconstruction. In the next subsection, we discuss the key principles for designing an FA block that efficiently manages computational resources while achieving state-of-the-art (SOTA) communication performance across various channel conditions.
Figure 3: Overview of FA block.
III-A Design Principles
Our design is inspired by a key observation in computer vision: not all data features contribute equally to task performance. Some features are highly informative, while others are less informative [32]. Hence, allocating more computational resources to informative features enhances data reconstruction, whereas assigning excessive computation to less informative features either provides negligible or limited performance gains [34, 35]. In other words, applying the same computational resource to both important and less important features, as in existing deepJSCC works [6, 7, 8], is an inefficient use of deepJSCC’s computational resources. Based on this insight, we propose an efficient FA block that processes features differently according to their importance and dynamically adjusts the number of important features assigned to a high computational load within a single neural network. Additionally, for the FA block to be SNR-adaptive within a single neural network, randomly selected SNR values are given for the sake of simplicity as in [7, 8, 13]. In the following steps, we present the design principles of our FA block as illustrated in Figure 3.111Note that the figure (and subsequent figures as well) illustrates the processing in the image domain to ease understanding, but the actual FA block operates in the feature domain.
  1. 1.
    SNR-Adaptive Feature Extraction: To adapt to varying channel SNRs, the FA block must incorporate random SNR information to the input feature $z_{{\textrm{in}}}$. Since feature distributions vary between FA blocks, each FA block requires an SNR feature that aligns with its own feature distribution to preserve the information in $z_{{\textrm{in}}}$. To achieve this efficiently, the SNR features should be extracted from the given SNR value with lightweight computation, ensuring it is tailored to each FA block.
  2. 2.
    Feature Classification: To allocate computational resources efficiently under a given computational budget, important and less important regions in the feature should be correctly identified. If the size of a feature is large, it should be split into several small-sized features. Based on the computational budget and the importance of each small-sized feature, small-sized features need to be properly classified into two categories: important and less important.
  3. 3.
    Preserving Information: Less important features may also contain some relevant information about the adjacent important ones. In Step $2$ of Figure 3, important features (upper) contain the most information about the flower. However, some less important features (lower) near the important ones also contain a small amount of information about the flower. Thus, important information in less important features needs to be incorporated into the near important ones to preserve important information during further processing.
  4. 4.
    Importance-Wise Feature Computation: Customized computations should be applied to classified features based on importance. Important features require intensive, high-capacity processing to enhance task performance, while less important features require efficient, lightweight computation to reduce computational overhead.
  5. 5.
    Unified Light-Weight Computation: Applying customized operations to classified features is efficient and effective for computer vision tasks [34]. However, in communication scenarios, both source coding’s compression and channel coding’s error correction must be considered together. From a source coding perspective, compressing redundant information as much as possible is crucial for description length. From a channel coding perspective, leveraging feature correlations to correct errors caused by channel noise is essential. Simply repeating customized operations for selected features does not address these two key aspects. To overcome this limitation, a unified lightweight computation for all features is required, one that compresses redundant information while also leveraging feature correlations after applying customized operations.
In the next subsection, we explain how we implement the design principles in the FA block.
III-B FA Block Details
Figure 4: (Step 1) Two-step SNR feature extraction for each FA block.
Two-Step SNR Feature Extraction: To implement an SNR-adaptive JSCC in a single model, FAJSCC employs a two-step SNR feature extraction process from random SNR, as shown in Figure 4. First, a randomly given SNR value is preprocessed by a multi-layer perceptron (MLP) at the FA group level. Then, the preprocessed SNR information is refined into a block-wise SNR feature through a fully connected (FC) layer for each FA block. Finally, each block-wise SNR feature is added to the FA input feature $z_{{\textrm{in}}}$ to incorporate SNR information, as shown in Figure 3.
Since the SNR features for FA blocks are preprocessed by a common MLP in each FA group and refined by a block-wise single FC layer, this two-step SNR information extraction approach introduces minimal computational overhead. Moreover, the specialized FC layers for each FA block enable refining SNR features to align with the input feature distributions of their respective FA blocks. These properties fulfill the key design principles of SNR feature extraction: efficiently extracting SNR features to incorporate SNR information into the FA input feature $z_{{\textrm{in}}}$ while preserving its original information. Based on this, FAJSCC can adapt to varying channel conditions, achieving improved performance compared to previous deepJSCCs. When the channel SNR is fixed and SNR-adaptive ability is no longer needed, this functional block is deleted; in this case, $z_{\textrm{in}}$ is directly fed into the next step.
Figure 5: (Step 2) Based on importance ratio $\gamma$ and patch-wise important scores, each feature patch is classified into important or less important features.
Feature Classification: To distinguish between important and less important regions in the feature, the input feature $z_{{\textrm{in}}}$ with $H_{z_{{\textrm{in}}}}\times W_{z_{{\textrm{in}}}}\times C_{z_{{\textrm{in}}}}$ dimension is split into $\frac{H_{z_{{\textrm{in}}}}W_{z_{{\textrm{in}}}}}{w^{2}}$ feature patches, each with $w\times w\times C_{z_{{\textrm{in}}}}$ dimension where $w\times w$ represents the patch resolution. Then, the FA block predicts the importance score of each patch using an importance predictor [35]. The importance predictor assigns importance scores ranging from 00 to $1$ by merging $w\times w$ patch resolution into $1\times 1$ resolution and applying MLP. Informally, the importance score reflects the efficiency of performance improvement when allocating higher computational resources to a given feature. In other words, patches that yield greater performance improvements when assigned more computational resources receive higher importance scores. As the feature classification is a discrete-output process, we need a special method to enable backpropagation. The Gumbel-Softmax trick enables the training, which will be detailed in Section III-C.
To maximize computational efficiency while adapting to varying computational budgets, the FA block classifies the top $\lfloor\frac{\gamma H_{z_{{\textrm{in}}}}W_{z_{{\textrm{in}}}}}{w^{2}}\rfloor$ feature patches with the highest importance scores as important, while treating the remaining patches as less important, as shown in Figure 5. The computational cost can be adjusted by setting the FA block’s importance ratio $\gamma$ to match the available computational budget. Since important features are selected in order of their importance scores, our FA block optimally balances performance and computational efficiency.
Figure 6: (Step 3) Detailed Procedures of feature preserving process.
Figure 7: (Step 3) The starting points of the green arrows indicate the locations of specific features, while the endpoints show where their related features are located. Together, these green arrows represent the offset, which defines the relative positions of the related features with respect to the given features. For simplicity, only the arrows for important features are visualized.
Preserving Information by Grid Sampling: To retain the relevant information of each important feature, our FA block employs grid sampling with an offset. The detailed feature preservation process is illustrated in Figure 6. First, the full feature, which contains both important and less important features, is used by the offset predictor [35] to estimate the offset. The offset is visualized in Figure 7, where the starting points of the green arrows indicate the locations of specific features, while the endpoints indicate the locations of their related features. These green arrows collectively represent the offset, which determines where the related features are located relative to the given features.
After predicting the offset, grid sampling extracts less important features related to important ones from the full feature with the offset. Finally, important features retain relevant information by incorporating their corresponding related features.
Importance-Wise Feature Computation: An intensive computation method for important features and a lightweight computation method for less important features should be considered for efficient importance-wise feature computation. To do this, our FA block employs self-attention [36, 37, 38] for intensive computation and spatial attention [39] for lightweight computation. The self-attention mechanism is highly effective for processing large information in various communication scenarios [41, 16, 8, 9, 10]. On the other hand, the spatial attention mechanism is highly efficient for feature refinement ability [39]. Thus, these two mechanisms are well-suited for customized importance-wise feature processing. Note that we use single-head self-attention, unlike multi-head self-attention used in SwinJSCC [8], to decrease computational resource and memory usage.
Unified Light-Weight Computation: Although importance-wise feature computation with different processing is verified as an efficient method in deep learning literature [34], relying solely on this method is not sufficient for communication tasks. In communication, leveraging the feature correlations is important for error correction. To assist this with a low computational burden, after importance-wise feature processing, our FA block employs depth-wise convolution operation [42] and a fully connected layer. On the encoder side, depth-wise convolution operation and a fully connected layer increase spatial-wise and channel-wise feature correlations. On the decoder side, these operations help correct feature errors caused by channel noise by utilizing the correlation of the features.
Another important aspect of successful communication is filtering out unimportant or noisy information. To achieve this with low computational burden, after depth-wise convolution operation and a fully connected layer, our FA block processes feature with a gated depth-wise convolution feed-forward network (GDF) [43]. On the encoder side, GDF compresses unimportant information to fit communication capacity. On the decoder side, GDF filters out highly corrupted information caused by communication noise.
III-C Training and Loss Function Design
During the test phase, patch-wise feature classification for importance-wise feature computation is sufficient for inference. However, this discrete classification prevents backpropagation during the training. To address this issue, the FA block employs hard Gumbel-Softmax sampling [44] during training. This approach replaces non-differentiable classified features with differentiable sampled features, enabling proper gradient flow. More specifically, the Gumbel-Softmax method samples important features with probabilities proportional to their importance scores. Thus, features that yield great performance improvements when assigned more computational resources should receive high-importance scores to be sampled as important features.

### v2

@@ III FAJSCC Framework
Our overall FAJSCC structure is illustrated in Figure 2. To facilitate deepJSCC training, standardization of the encoder maps pixel values of an image from $[0,255]$ to [1,1][-1,1] while destandardization of the decoder reverses this process. Power normalization of the encoder rescales features to fit the power constraint $P=1$. Following previous works [40, 16, 8, 10], our FAJSCC uses patch embedding to refine the image at the feature level, patch merging for down-sampling, and patch division for up-sampling. Based on the feature sizes, both the encoder and decoder consist of multiple stages, indexed by $i\in[1:L]$, where $L$ is the total number of stages. The patch merging layer of the $i$-th stage transforms the feature dimension from $\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times C_{i-1}$ to $\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{i}$. Conversely, the patch division layer in the $i$-th stage transforms the feature dimensions from $\frac{H}{2^{L+1-i}}\times\frac{W}{2^{L+1-i}}\times C_{L+1-i}$ to $\frac{H}{2^{L-i}}\times\frac{W}{2^{L-i}}\times C_{L-i}$. Here, $C_{0}$ is the number of RGB color channels, i.e., $C_{0}=3$. The encoder’s fully connected (FC) layer scales input feature channel dimensions to fit the available bandwidth, while the decoder’s FC layer adjusts input feature channel dimensions to align the feature channel sizes of the decoder’s first stage.
Figure 3: Dimension specialized operations of the proposed FA block.
In each stage, the feature importance-aware group (FA group), consisting of multiple feature importance-aware blocks (FA blocks), performs the core encoding and decoding processes of FAJSCC. To enhance communication accuracy while reducing computational resources, our FA block integrates axis-dimension specialized computation and selective deformable self-attention within an attention family tree. Figure 3 depicts the overall procedure of the proposed dimension-specialized operations. The axis-dimension specialized computation refines features separately along the channel and spatial axes, reducing computational cost while representing features effectively. It corresponds to the main computational path except for the selective deformable self-attention module. The selective deformable self-attention applies self-attention only to the most important features by adaptively adjusting attention regions according to the input feature relations and the selected windows. Furthermore, the proposed attention family tree provides the attention features required for both axis-dimension specialized computation and selective deformable self-attention. This is achieved with highly efficient computation by eliminating redundant calculations based on output feature relationships. The following subsections present our proposed methods in detail.
III-A Axis-Dimension Specialized Computation
The standard convolution operation in deep learning layers processes features simultaneously along the height, width, and channel axes. While this joint processing is effective for learning correlations across all dimensions, it incurs a high computational complexity of $\mathcal{O}(H_{{\textrm{out}}}W_{{\textrm{out}}}C_{{\textrm{out}}}C_{{\textrm{in}}}s^{2})$. Here, $H_{{\textrm{out}}},W_{{\textrm{out}}},C_{{\textrm{out}}},C_{{\textrm{in}}}$ denote the output feature’s height, width, channels, and input feature’s channels, respectively, and $s$ represents the kernel size. This complexity can be reduced without sacrificing representational ability by decomposing the operation into two axis-specific convolutions: depthwise and pointwise. Depthwise convolution (dwConv) processes the spatial dimensions (height and width) independently within each channel, whereas pointwise convolution (pConv) operates only along the channel axis. Together, they capture spatial and channel correlations with complexities of $\mathcal{O}(H_{{\textrm{out}}}W_{{\textrm{out}}}C_{{\textrm{out}}}s^{2})$ and $\mathcal{O}(H_{{\textrm{out}}}W_{{\textrm{out}}}C_{{\textrm{out}}}C_{{\textrm{in}}})$, respectively, thereby reducing overall computational cost while achieving strong feature processing ability.
In addition to employing depthwise and pointwise convolutions, we apply a spatial attention operation (multiply input feature with spatial attention $\mathbf{z}_{sa}\in[0,1]^{H_{{\textrm{in}}}\times W_{{\textrm{in}}}\times 1}$) before the depthwise convolution and a channel attention operation (multiply input feature with channel attention $\mathbf{z}_{ca}\in[0,1]^{1\times 1\times C_{{\textrm{in}}}}$) before the pointwise convolution. Spatial attention emphasizes important features along the spatial axes, while channel attention emphasizes important features along the channel axes [38]. These attentions enhance the depthwise and pointwise convolutions by informing which features are important. Moreover, these lightweight attention operations incur only $\mathcal{O}(H_{{\textrm{in}}}W_{{\textrm{in}}}C_{{\textrm{in}}})$ complexity, imposing minimal computational overhead. Using depthwise and pointwise convolutions has also been adopted in recent deepJSCC research [32]. However, the application of lightweight attentions to enhance these operations is a novel idea, and its effectiveness is further validated through the ablation study in Section IV-C, in comparison with using depthwise and pointwise convolutions alone.
III-B Selective Deformable Self-Attention
Although the above axis-dimension specialized operations effectively refine features along each axis, relying solely on them may overlook relations among complex features. To efficiently capture these relations, we propose the deformable self-attention mechanism. The detailed self-attention operation for $\mathbf{Z}_{{\textrm{in}}}\in\mathbb{R}^{H_{{\textrm{in}}}W_{{\textrm{in}}}\times C_{{\textrm{in}}}}$ is
$\mathbf{Z}_{{\textrm{out}}}=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\intercal}}{\sqrt{C_{{\textrm{in}}}}}\right)\mathbf{V},$
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{H_{{\textrm{in}}}W_{{\textrm{in}}}\times C_{{\textrm{in}}}}$ are the query, key, and value matrices obtained from $\mathbf{Z}_{{\textrm{in}}}$ through learnable linear projections, respectively. This full self-attention has been shown to be highly effective in semantic communication [33, 37], as it captures dependencies among complex features through feature similarity measurement with $\mathbf{Q}\mathbf{K}^{\intercal}$. However, its computational complexity $\mathcal{O}\left(4H_{{\textrm{in}}}W_{{\textrm{in}}}C_{{\textrm{in}}}^{2}+2(H_{{\textrm{in}}}W_{{\textrm{in}}})^{2}C_{{\textrm{in}}}\right)$ becomes prohibitively large as image resolution increases. To alleviate this computational burden, many semantic communication systems adopt window-based self-attention [40, 16, 8, 9, 41, 10]. Since this method applies self-attention only within fixed local windows, the computational complexity is reduced to $\mathcal{O}\left(\frac{H_{{\textrm{in}}}W_{{\textrm{in}}}}{n}\left(4nC_{{\textrm{in}}}^{2}+2n^{2}C_{{\textrm{in}}}\right)\right)$ where $n$ is the number of features in the given window and $n\ll H_{{\textrm{in}}}W_{{\textrm{in}}}$ in most cases. While conventional window self-attention enhances computational efficiency, this method has two major limitations: It cannot capture highly related features across different windows, and analyzing features of every window is computationally expensive. We solve these problems by proposing selective deformable self-attention.
Figure 4: Visualization of feature deformation.
To capture strongly related features across different windows, we apply feature deformation before performing self-attention, which relocates related regions. The detailed process of feature deformation for the $i,j,c$-th height, width, and channel is as follows:
$(u,v)=(i,j)+\mathbf{z}_{{\textrm{offset}}}(i,j),\quad i_{0}=\lfloor u\rfloor,\;i_{1}=i_{0}+1,$
$j_{0}=\lfloor v\rfloor,\;j_{1}=j_{0}+1,\quad\alpha=u-i_{0},\;\;\beta=v-j_{0},$
$z_{{\textrm{out}}}^{c}(i,j)=\begin{bmatrix}1-\alpha&\alpha\end{bmatrix}\begin{bmatrix}z_{{\textrm{in}}}^{c}(i_{0},j_{0})&z_{{\textrm{in}}}^{c}(i_{0},j_{1})\\ z_{{\textrm{in}}}^{c}(i_{1},j_{0})&z_{{\textrm{in}}}^{c}(i_{1},j_{1})\end{bmatrix}\begin{bmatrix}1-\beta\\ \beta\end{bmatrix}$.
Here, $(u,v)$ denotes the deformed sampling location predicted from the original coordinate $(i,j)$ and the predicted offset $\mathbf{z}_{{\textrm{offset}}}\in\mathbb{R}^{H_{{\textrm{in}}}\times W_{{\textrm{in}}}\times 2}$ from the input feature relations. The four neighboring integer grid points $i_{0},i_{1},j_{0},j_{1}$ represent the surrounding pixels (or features) that are closest to $(u,v)$. The interpolation weights $\alpha$ and $\beta$ correspond to the fractional distances between $(u,v)$ and the integer coordinates along the height and width axes, respectively. The equation for $z_{{\textrm{out}}}^{c}(i,j)$ performs bilinear interpolation: When $(u,v)$ coincides with an integer grid point, it directly samples the corresponding feature value, whereas when $(u,v)$ lies between grid points, it computes a weighted combination of the four surrounding features, thereby interpolating the feature map at the non-integer position. This formulation ensures that both cases (direct alignment with existing feature coordinates and interpolation at in-between positions) are handled within a single equation, allowing subsequent self-attention to adaptively operate on adjusted regions instead of being limited to fixed grids. This effect is illustrated in Figure 4.111Note that the figure (and subsequent figures as well) illustrates the deformation in the image domain to ease understanding, but the actual deformation operates in the feature domain. The thick black borders in this figure indicate the features selected for self-attention. Applying self-attention within fixed-window areas may overlook feature correlations beyond them. In contrast, feature deformation enables self-attention to operate on adaptively adjusted regions, thereby capturing correlations that go beyond the fixed windows. We call this method deformable (window) self-attention.
The idea of feature deformation originates from deformable convolution [46, 47], which has also been successfully applied in recent semantic communication systems [48]. In addition, feature deformation for full self-attention has been explored in computer vision tasks [49], where the entire input feature space is deformed. However, such global deformation is not suitable for the selective enhancement framework that we introduce in the next paragraph, because it operates over the entire feature set and therefore cannot support selective refinement. In contrast, our method confines the deformation to features within a given window, enabling more targeted and efficient enhancement. Furthermore, to the best of our knowledge, feature deformation has not previously been applied to self-attention for constructing a novel deepJSCC architecture.
Deformable self-attention effectively enhances the capability of standard self-attention. Building on this, we propose to applying deformable self-attention selectively to improve computational efficiency. Previous studies in neural source coding have shown that not all features carry the same amount of information: Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [42]. Moreover, it has been demonstrated that simple images do not require high computational cost to achieve high performance in super-resolution tasks [43, 44]. Motivated by these observations, we apply deformable self-attention only to features in the most important windows, identified based on their contribution to image transmission performance. By restricting self-attention to these windows, the computational cost is significantly reduced. This selective enhancement strategy greatly reduces computational cost compared with applying deformable self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to features in all windows. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity.
Figure 5: Procedures of selective deformable self-attention.
Figure 5 gives the detailed procedures of our proposed selective deformable self-attention. Firstly, based on the window importance $\mathbf{z}_{{\textrm{imp}}}\in[0,1]^{\frac{H_{{\textrm{in}}}}{\sqrt{n}}\times\frac{W_{{\textrm{in}}}}{\sqrt{n}}}$, windows are selected whether to be enhanced via deformable self-attention or not. The number of selected windows is determined by the importance ratio $\gamma$. When $\gamma=1$, features in all windows are processed via deformable self-attention, whereas $\gamma=0$ means none are processed. For $0<\gamma<1$, the features in the top $\gamma$ fraction windows, ranked by their measured importance, are selected and processed through deformable self-attention, which results in $\mathcal{O}\left(\frac{\gamma H_{{\textrm{in}}}W_{{\textrm{in}}}}{n}\left(4nC_{{\textrm{in}}}^{2}+2n^{2}C_{{\textrm{in}}}\right)\right)$ computational complexity. Secondly, the feature deformation process of Figure 4 is applied to the selected windows to adjust self-attention regions based on the input feature relations. Lastly, self-attention is applied to the selected adjusted areas and restore the original input feature shape with processed selected features and unselected features.
Unlike previous lightweight deepJSCC methods that reduce or remove features [36, 33, 37], which often degrade performance, FAJSCC selectively enlarges and refines important features, achieving higher performance while reducing computational cost. Moreover, by adjusting the importance ratio, FAJSCC decouples computational complexity from the model architecture, allowing first independent evaluation of the impact of encoder and decoder complexity on performance. The prior resource-adjustable methods [33, 37] cannot achieve this, as the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [33, 37]. Experiments varying encoder and decoder resources independently reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. Detailed results are shown in Section IV-B.
III-C Attention Family Tree
The proposed axis-dimension specialized operation and the selective deformable self-attention are both highly efficient and impose minimal computational overhead. However, we observe that naïvely computing the attention features, offsets, and window importance from scratch negates the efficiency gains of our methods; thus, we propose a new attention family tree that jointly extracts attention features and obtains offset and feature importance at once by eliminating overlapping calculations based on relations among output features to minimize computational resource overhead.
Note that the offset used in the deformation is obtained from the spatial-wise feature information, and spatial attention is obtained from the spatial-wise feature importance. The spatial-wise feature importance can be obtained from the spatial-wise feature information; thus, reusing the spatial-wise feature information to get spatial-wise feature importance can eliminate overlapping pre-processing calculations. Adding to this, spatial-wise feature importance can be directly used to identify which window is important, as well as obtain the spatial attention feature. Thus, the attention, offset, and window importance can all be obtained efficiently without a large computational burden. In ablation studies in Section IV-C, we show that our attention family tree shows better or similar performance while reducing computational burden compared to obtaining these attention features separately. Moreover, we also show that the efficiency gain comes from appropriately leveraging the relations of offset, attention feature, and window importance.
III-D Training and Loss Function Design
Unlike in the test phase where simply selecting the top $\gamma$ fraction of windows suffices for selective deformable self-attention, such discrete selection prevents backpropagation during training. To address this issue, we employ Gumbel-Softmax sampling [50] during training. This approach replaces a non-differentiable selection process with a differentiable sampling process, enabling proper gradient flow. More specifically, we sample important windows with probabilities proportional to window importance $\mathbf{z}_{{\textrm{imp}}}$. By defining the probability that $(i,j)$-th coordinate window is sampled as important window $\pi^{1}(i,j):=z_{{\textrm{imp}}}(i,j)$, and the probability that $(i,j)$-th coordinate window is not sampled as important window $\pi^{2}(i,j):=1-z_{{\textrm{imp}}}(i,j)$, the soft sampling result $\mathbf{z}_{{\textrm{soft}}}$ and the hard sampling result $\mathbf{z}_{{\textrm{hard}}}$ are as follows:
$z^{l}_{{\textrm{soft}}}(i,j)$ $=\frac{\exp{\left(\log\pi^{l}(i,j)+g^{l}(i,j)\right)/\tau}}{\sum_{m=1}^{2}\exp{\left(\log\pi^{m}(i,j)+g^{m}(i,j)\right)/\tau}},$
$\mathbf{z}_{{\textrm{hard}}}(i,j)$ $={\textrm{one\_hot}}\left(\operatorname*{arg\,max}_{l}\left[\log\pi^{l}(i,j)+g^{l}(i,j)\right]\right).$
Here, $g^{l}(i,j)$ is i.i.d. sample from $\text{Gumbel}(0,1)$, [^2] This sampling can be done with $t^{l}(i,j)\simUniform$$(0,1)$, and $g^{l}(i,j)=-\log(-\log(t^{l}(i,j)))$. $\tau$ is a hyperparameter, and ${\textrm{one\_hot}}(\cdot)$ is a one-hot vector whose entry is $1$ at the selected index and 0 elsewhere.
The hard sampling result $\mathbf{z}_{{\textrm{hard}}}$ follows the exact true distribution with $\mathbf{\pi}^{1},\mathbf{\pi}^{2}$ probabilities, but this prevents backpropagation. In contrast, the soft sampling result $\mathbf{z}_{{\textrm{soft}}}$ enables backpropagation, but it deviates to some extent from the true distribution, except when $\tau=0$, i.e., $\mathbf{z}_{{\textrm{soft}}}=\mathbf{z}_{{\textrm{hard}}}$. To leverage the advantages of $\mathbf{z}_{{\textrm{soft}}},\mathbf{z}_{{\textrm{hard}}}$ both, we use the straight-through important window sampling result $\mathbf{z}_{{\textrm{ST}}}$ as follows:
$\mathbf{z}_{{\textrm{ST}}}=\mathbf{z}^{1}_{{\textrm{hard}}}+\mathbf{z}^{1}_{{\textrm{soft}}}-{\textrm{stop\_grad}}(\mathbf{z}^{1}_{{\textrm{soft}}}).$
Here, ${\textrm{stop\_grad}}(\cdot)$ indicates the stop-gradient operation, which behaves as the identity function in the forward propagation and outputs zero gradient in the backward propagation. The use of $\mathbf{z}_{{\textrm{ST}}}$ enables discrete sampling during the forward process while preserving gradient flow at the backward process during training. In selective deformable self-attention, we use $\mathbf{z}_{{\textrm{ST}}}$ to model the discrete sampling process and enable gradient flow as follows:
$\mathbf{z}_{{\textrm{out}}}={\textrm{self\_attention}}({\textrm{deformation}}(\mathbf{z}_{{\textrm{ST}}}\odot_{\textrm{w}}\mathbf{z}_{{\textrm{in}}})).$
Here, $\odot_{\textrm{w}}$ denotes window-wise multiplication that multiplies $z_{{\textrm{ST}}}(i,j)$ to the features in the $(i,j)$-th window of $\mathbf{z}_{{\textrm{in}}}$. Due to gradient flow via $\mathbf{z}^{1}_{{\textrm{soft}}}$ term of $\mathbf{z}_{{\textrm{ST}}}$, the window importance head in Figure 3 can be optimized with respect to a given loss function during training. Recall that this Gumbel-Softmax method is only used in the training phase. In the test phase, since gradient flow is not required, we do not use $\mathbf{z}_{{\textrm{ST}}}$ and simply select important windows in $\mathbf{z}_{{\textrm{in}}}$ based on $\mathbf{z}_{{\textrm{imp}}}$ as detailed in Section III-B.